Circular No. 7606

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 G1 An apparently asteroidal object discovered on CCD images taken with

the LONEOS 0.59-m Schmidt telescope (discovery observation given below) and posted on the NEO Confirmation Page has been found by other astrometric observers to be cometary on their CCD images. The object was reported as being diffuse by J. Tichá, M. Tichý, and P. Jelínek at Kleť (Apr. 1.9 and 2.9 UT; 9′′ coma on Apr. 2.9) and by C. E. Lopez and M. R. Cesco at El Leoncito (Apr. 2.2), and as having a 10′′ coma by M. Busch and S. Kluegl at Heppenheim, Germany (Apr. 1.9) and by J. Broughton, Reedy Creek, Queensland (Apr. 2.5). Additional astrometry and very uncertain parabolic orbital elements are given on MPEC 2001-G03.

2001 UT  α_{2000}  δ_{2000}  m_{2}
Apr. 1.19878  11^{h}31^{m}41.^{s}97  −21^{o}01′28′′.7  17.4

COMETS C/1996 L1 AND C/1997 M5 (SOHO) Further to IAUC 7601, D. Biesecker reports his measurements for two

Kreutz sungrazing comets found on LASCO C2 coronagraph images by X. Leprette at the SOHO website. The reduced observations and orbital elements by B. G. Marsden appear on MPEC 2001-F52.

Comet  Date  UT  α_{2000}  δ_{2000}
C/1996 L1  1996 June 10.913  5^{h}14.^{m}8  +21^{o}22′
C/1997 M5  1997 June 19.725  5 49.0  +21 46

SUPERNOVA 2001aj IN UGC 10243 D. Weisz and W. D. Li, University of California at Berkeley, report the

discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.0) on an unfiltered CCD image taken on Mar. 30.5 UT (and confirmed at ∼ the same brightness on an earlier image taken on Mar. 10.5) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2001aj is located at α = 16^{h}10^{m}30^{s}.23, δ = +19^{o}57′01′′.2 (equinox 2000.0), which is 9′′.0 east and 13′′.5 south of the nucleus of UGC 10243. A KAIT image taken on 2000 Aug. 19.3 showed nothing at this position (limiting mag ∼ 19.0). V. Pozzoli, Bareggio, Italy, reports an independent discovery of SN 2001aj at mag 18.1 on an unfiltered CCD image taken on Apr. 1.066, adding that nothing is present at this location on Digital Sky Survey images.

2001 April 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green